OpenStudy (anonymous):

solve for x 3x^2 + 2x-7=0

OpenStudy (anonymous):

Do you know about Quadratic Formula??

OpenStudy (anonymous):

yess i know how to do it

OpenStudy (anonymous):

By comparing it with \(ax^2 + bx + c = 0\) tell me what is a, b and c??

OpenStudy (anonymous):

Find Discriminant then: \(D = b^2 - 4ac\)

OpenStudy (anonymous):

Able in doing or not? @katiebugg

OpenStudy (jiteshmeghwal9):

given: a=3 b=2& c=-7 now plug this constants in the formula\[-b \pm \sqrt{-b^2-4ac} \over 2a\]

OpenStudy (jiteshmeghwal9):

sorry it's \[-b \pm \sqrt{b^2-4ac} \over 2a\]

OpenStudy (jiteshmeghwal9):

able to do it @katiebugg ???

OpenStudy (anonymous):

Firstly find D..

OpenStudy (anonymous):

mmk i no all the quadratci stuff my head just hurts and i cant doo itt at the moment so couldu u help

OpenStudy (anonymous):

Common excuse.. So, I can't help you if you will not do it yourself.. We can help you but cannot solve it for you..

OpenStudy (anonymous):

sorry doing like 15 others too

OpenStudy (anonymous):

If you know the quadratic formula then I think you are very well to do them..